Grade 11Mechanics

if vector A=3i+4j &vaector B=7i+24j.find the vector
having same magnitude as vector B & parallel to vector A

varenya joshi
10 Years agoGrade 11

4 Answers

Vikas TU
10 Years ago
Laet the new vector be C =ai + bj
a^2 + b^2 = 625         .....(1)
a/3 = b/4
a = (3/4)b      .......(2)
soving eqns. (1)  and  (2) 
a = 15 and a = -15
b = 20 and b =- 20
C = 15i + 20j
and C = 15i – 20j
saini
10 Years ago
a vector which is parallel to 3i+4j can be written as, let D is the vector then D=a(3i+4j)  where (a) is some constant as magnitude of B vector is 25 , and acc. to question the magnitude of D= magnitude of B, this implies that as magnitude of D is 5a so 5a = |25| and we get  a=+5 ,-5 and required vector is D=15i+20j or -(15i+20j) as the question say that the required vector is parallel to A vector so required answer should be D=15i+20j because negative sign says that it is antiparallel
habtamu kiflu
8 Years ago
 
if vector A=3i+4j &vaector B=7i+24j.find the vector
having same magnitude as vector B & parallel to vector A
Sai Disha
8 Years ago
Le